# The Memorization Problem: Can We Trust LLMs’ Economic Forecasts?

### v1

@@ 4.5 Portfolio Stock Price Recall by Size
To broaden our analysis of GPT-4o’s memorization, we also examine its recall of end-of-month closing prices for portfolios of stocks grouped by market capitalization—Small, Medium, and Large—covering January 1990 to September 2023, all within the model’s training cutoff of October 2023. Stocks are divided into market cap terciles, with 10 stocks randomly sampled from each tercile and resampled annually to account for size changes, using stock price data from CRSP. We test recall without context, with the previous two months’ prices provided, and over a recent 10-year period (2014–2023) to assess recency effects. Performance metrics, including Mean Percent Error, Mean Absolute Percent Error, and Directional Accuracy (correct direction of change relative to the prior month), are reported in Table 5. We also plot the actual vs estimated values in Figure 7.
Table 5: Evaluation Metrics for Portfolio of Stocks by Size
@@
This table reports a set of evaluation metrics for portfolios of stocks grouped by size: Small, Medium, and Large. We divide stocks into terciles by market cap (using the NYSE stocks only to calculate cutoffs) and sample 10 stocks from each tercile, resampling each year. We then ask the LLM to recall end of month closing prices for each stock. Mean Percent Error (MPE), Mean Absolute Percent Error (MAPE), and Directional Accuracy Change are reported in percentage points (0.78 means 0.78%). MPE is calculated by taking the average of the percent error $(PredictedPrice-ActualPrice)/ActualPrice$. MAPE is calculated by taking the average of the absolute value of the percent error. Directional Accuracy Change is the proportion of predictions that went in the correct direction (up or down) relative to the previous month. Confidence Calibration is the correlation between the LLM’s confidence level (on a scale from 0 to 100) and the mean absolute percent error. Num Obs is the number of observations used in the evaluation. Start Date and End Date indicate the period over which the metrics were computed. Refusals represent the number of instances in which the model withheld a prediction by either answering ”null” or 0. Results are provided for the full sample period, a recent period covering the past 10 years, and With Context, in which the previous two months’ closing prices are provided to the model.
MPE (%) MAPE (%) Directional Accuracy Change (%) Confidence Calibration Start Date End Date Num Obs Refusals
Small -3.50 18.35 46.58 -0.26 01/31/1991 09/29/2023 2539 1441
Medium -2.65 15.94 51.39 -0.31 12/31/1990 09/29/2023 2859 1159
Large -4.42 11.36 60.52 -0.37 01/31/1990 09/29/2023 3188 837
Recent Period: Past 10 years
Small 0.32 9.79 52.54 -0.04 01/31/2014 09/29/2023 1145 4
Medium 0.77 5.94 64.84 -0.00 01/31/2014 09/29/2023 1158 3
Large 0.57 3.57 77.03 -0.07 01/31/2014 09/29/2023 1158 1
With Context
Small 0.52 8.84 74.57 -0.03 01/31/1990 09/29/2023 3856 118
Medium 1.02 8.20 73.89 -0.04 01/31/1990 09/29/2023 3883 111
Large 1.11 6.59 74.70 -0.10 01/31/1990 09/29/2023 3803 158
@@
The results show weaker recall compared to the Magnificent 7, with accuracy improving for larger stocks and recent periods. Without context, Mean Absolute Percent Errors are high: 18.35% for Small, 15.94% for Medium, and 11.36% for Large stocks, with Directional Accuracy ranging from 46.58% to 60.52%. The high refusal rates (1441 for Small, 837 for Large) suggest uncertainty for less prominent stocks. Over the recent 10-year period, errors decrease significantly—9.79% for Small, 5.94% for Medium, and 3.57% for Large—with Directional Accuracy rising to 52.54%–77.03% and near-zero refusals, echoing the recency effect seen in macroeconomic levels. With context, errors further improve for Directional Accuracy with all three sizes reaching at least 73%. Larger stocks consistently show better recall, likely due to greater data prominence, aligning with the Magnificent 7’s stronger performance for high-profile securities.
Figure 7: Recall of exact numerical levels of closing prices for other stocks.
This figure shows the LLM’s estimated closing prices for randomly selected small, mid, and large cap stocks compared to the actual closing prices. Panels A, C, and E graph the actual values against the estimated values. Panels B, D, and F show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%). The price plotted is the equal-weighted average price of the small, mid and large cap groups. For the large stock portfolio, ticker BRK was excluded for ease of plotting due to the extreme jump in average price with its inclusion.
@@
These findings complement our Magnificent 7 results, revealing that GPT-4o’s memorization weakens for less prominent stocks, particularly smaller ones, and is strongest for recent, larger-cap data. The high errors and refusals for small stocks contrast with the precision for META or GOOGL, suggesting memorization is skewed toward widely covered securities, similar to the robust recall of market indices and macroeconomic rates. The recency and context effect parallel improvements seen in prior tests, but the lower directional accuracy with context indicates limits in capturing trends for diverse portfolios. This selective memorization underscores the risk of relying on LLMs for historical stock analysis, as apparent forecasting accuracy may stem from memorized prices of prominent stocks rather than broad predictive insight, reinforcing the need for post-cutoff evaluations.
@@ 4.6 Can LLMs Follow “Fake” Knowledge Cutoff Prompts?
To assess whether GPT-4o can adhere to instructions not to use information beyond a specified cutoff, we test its performance in predicting U.S. real GDP growth rates using data from the FRED database, spanning March 1990 to June 2023. We design three prompting conditions using an artificial cutoff date of December 2010: (i) one where both system and user messages explicitly restrict knowledge to pre-2010 data, (ii) one where only the system prompt imposes this fake cutoff, and (iii) one where only the user prompt imposes this fake cutoff. The task required predicting monthly year-over-year GDP growth, with data split into pre-cutoff (1990–2010) and post-cutoff (2011 onward) periods to evaluate compliance with the constraint. Performance metrics include Mean Percent Error, Mean Absolute Percent Error, Directional Accuracy (correctly identifying up/down changes), and refusal counts, reported in Table 6, building on our prior findings by directly probing the model’s ability to isolate historical forecasting from memorized data.
@@
When both system and user prompts enforce the pre-2010 cutoff, GPT-4o performs well on pre-cutoff data, with a Mean Absolute Percent Error of 0.27% and Directional Accuracy of 95.18%, but struggles post-cutoff, showing a higher error (0.59%) and lower accuracy (92.31%), with 38 refusals out of 51 post-cutoff observations. This high refusal rate and reduced performance suggest the model partially respects the constraint, limiting its reliance on memorized post-2010 data. We also find that the results are similar when only the user prompt specifies the 2010 cutoff. In contrast, when only the system prompt specifies the cutoff, pre-cutoff accuracy improves slightly (0.11% error, 98.80% accuracy). Still, post-cutoff performance remains implausibly strong (0.05% error, 96.00% accuracy) with only one refusal across 50 observations. This minimal drop in accuracy post-cutoff indicates that without reinforced user instructions, the model likely accesses memorized data, undermining the cutoff constraint.
These results connect to our earlier findings on macroeconomic indicators, where high pre-cutoff accuracy reflected memorization. The strong post-cutoff performance without user prompt reinforcement mirrors the suspiciously high accuracy seen in other tests when constraints were not strictly enforced, suggesting that GPT-4o defaults to using its full knowledge unless explicitly and repeatedly directed otherwise. The high refusal rate with dual prompts aligns with weaker recall for less prominent data, as seen in small-cap stocks, indicating partial compliance but not complete isolation from memorized information. This failure to fully respect cutoff instructions reinforces the challenge of using LLMs for historical forecasting, as their outputs may subtly incorporate memorized data, necessitating post-cutoff evaluations to ensure genuine predictive ability.
Table 6: Knowledge Pool Cutoff
@@
This table reports GPT-4o’s performance on U.S. real GDP growth predictions evaluated on data from the Philadelphia Fed’s Real-Time Data Set. We evaluate model accuracy under different knowledge cutoff constraints: one where both system and user prompts reinforce the knowledge cutoff (pre-2010 only), another where only the system prompt specifies the cutoff, and one where only the user prompt specifies the cutoff. The task involves predicting the monthly year-over-year GDP growth rate, with test data split into pre-cutoff (1990–2010) and post-cutoff (2011 onward) periods to assess whether the model respects the stated cutoff. Metrics include Mean Error (ME), Mean Absolute Error (MAE), Directional Accuracy (percentage of guesses correctly above a threshold of 2.5%), Directional Accuracy Change (percentage of correct up/down changes), Confidence Calibration (correlation between the LLM’s confidence level and the MAPE), total observations, and refusal counts. The results indicate that explicitly instructing the model not to use post-2010 data yields higher refusal rates and weaker post-cutoff performance, consistent with adherence to the knowledge constraint.
@@
ME (%) MAE (%) Directional Accuracy (%) Directional Accuracy Change (%) Confidence Calibration Start Date End Date Num Obs Refusals
@@
Post fake-cutoff 0.58 0.59 69.23 75.00 0.51 03/01/2011 12/01/2016 13 38
@@
Post fake-cutoff -0.01 0.05 98.00 100.00 -0.24 03/01/2011 06/01/2023 50 1
@@
Post fake-cutoff 0.60 0.61 66.67 70.59 -0.15 03/01/2011 06/01/2020 18 32
@@ 4.7 Testing Masking Effectiveness
To evaluate whether masking can prevent GPT-4o from accessing memorized information, we examine its ability to identify the firm, year, and quarter of anonymized earnings conference call transcripts from Capital IQ, covering July 2006 to December 2021, all within the model’s training cutoff of October 2023. Transcripts were anonymized using the entity neutering approach of Engelberg et al. (2025), removing identifying details like company names and dates. We focused on the Magnificent 7 stocks (AAPL, META, MSFT, GOOG, NVDA, TSLA, AMZN) and portfolios grouped by market cap terciles (Large, Medium, Small), measuring performance through Mean Years Difference, Mean Absolute Years Difference, and accuracy in identifying the exact year, quarter and year, and firm. Results, reported in Table 7, extend our prior findings by probing whether masking mitigates the memorization seen in stock price and macroeconomic data recall.
Table 7: Anonymous Conference Calls
This table reports GPT-4o’s performance in identifying the correct firm, year, and quarter of anonymized earnings conference call transcripts obtained from Capital IQ. For each firm, we report the mean and mean absolute difference between the model’s predicted and actual transcript years, as well as the percentage of calls for which the model correctly identified the exact year, the exact quarter and year, and the correct firm. Results are grouped by firm and by market capitalization terciles (Large, Mid, Small). Terciles are formed using NYSE cutoffs.
@@
AAPL -0.03 0.06 95.24 92.06 100.00 63
META 0.21 0.26 73.68 2.63 100.00 38
MSFT -0.73 0.85 61.29 1.61 100.00 62
GOOG -1.29 1.83 57.14 7.94 90.48 63
NVDA -1.67 1.94 47.06 1.96 92.16 51
TSLA -4.11 4.20 33.33 4.44 82.22 45
AMZN -2.23 2.42 23.44 4.69 82.81 64
@@
Large -1.75 2.12 42.05 9.09 65.53 528
Mid -2.08 2.57 38.49 8.11 40.38 530
Small -1.69 2.44 31.67 6.43 32.14 420
@@
GPT-4o demonstrates a remarkable ability to deanonymize transcripts, particularly for prominent firms. For AAPL, the model achieves 100% firm accuracy, 95.24% year accuracy, and 92.06% quarter-and-year accuracy, with a Mean Absolute Years Difference of just 0.06 years. META and MSFT also show perfect firm identification (100%), though year accuracy drops to 73.68% and 61.29%, respectively, and quarter-and-year accuracy is low (2.63% and 1.61%). Performance weakens for GOOG (90.48% firm accuracy), NVDA (92.16%), TSLA (82.22%), and AMZN (82.81%), with year accuracy ranging from 23.44% to 57.14% and Mean Absolute Years Differences rising to 1.83–4.20 years. Across market cap terciles, firm accuracy declines from 65.53% for Large to 32.14% for Small stocks, with year accuracy dropping from 42.05% for Large to 31.67% for Small stocks, indicating stronger memorization for larger, more prominent firms, consistent with our portfolio stock findings.

### v2

@@ 5.4.2 Individual Stocks in the Cross Section
To broaden our analysis of GPT-4o’s memorization, we also examine its recall of end-of-month closing prices for individual stocks grouped by market capitalization covering January 1990 to December 2024. Stocks are divided into market capitalization quintiles with size breakpoints calculated using NYSE stocks only. We randomly sample 50 stocks from each quintile and resampled annually to account for size changes, using stock price data from CRSP. We test recall without context, with the previous two months’ prices provided, and over the recent 10-year pre-cutoff period to assess recency effects. Performance metrics, including Mean Percent Error, Mean Absolute Percent Error, and Directional Accuracy (correct direction of change relative to the prior month), are reported in Table 5. We also plot the actual vs estimated values in Figure 7.
Table 5: Evaluation Metrics for Individual Stocks by Size Quintile
@@
This table reports a set of evaluation metrics for portfolios of stocks grouped by size quintiles. We divide stocks into quintiles by market cap (using the NYSE stocks only to calculate cutoffs) and sample 50 stocks from each quintile, resampling each year. We then ask the LLM to recall end of month closing prices for each stock. Mean Percent Error (MPE), Mean Absolute Percent Error (MAPE), and Directional Accuracy are reported in percentage points (0.78 means 0.78%). MPE is calculated by taking the average of the percent error (PredictedPriceActualPrice)/ActualPrice(PredictedPrice-ActualPrice)/ActualPrice. MAPE is calculated by taking the average of the absolute value of the percent error. Directional Accuracy is the proportion of predictions that went in the correct direction (up or down) relative to the previous month. Confidence Calibration is the correlation between the LLM’s confidence level (on a scale from 0 to 100) and the mean absolute percent error. Num Obs is the number of observations used in the evaluation. The panel labels and subheadings indicate the period over which the metrics were computed. Refusals represent the number of instances in which the model withheld a prediction by either answering ”null” or 0. Results are provided for the full sample period, a recent period covering the past 10 years, and With Context, in which the previous two months’ closing prices are provided to the model.
Panel A: No Context
Quintile MPE (%) MAPE (%) Directional Accuracy (%) Confidence Calibration Num Obs Refusals
Pre-cutoff, 01/1990 to 09/2023
1 1.20 24.70 42.66 -0.10 11,925 6828
2 -0.55 19.17 48.78 -0.13 12,262 6394
3 -3.04 17.06 51.02 -0.32 13,077 5469
4 -3.36 14.18 53.64 -0.22 13,593 4908
5 -4.55 11.29 61.11 -0.26 14,816 3576
Recent Period, 01/2014 to 09/2023
1 4.24 15.24 48.62 -0.03 5,415 96
2 2.71 10.18 58.59 -0.04 5,448 26
3 1.53 6.45 63.61 -0.02 5,403 20
4 0.54 4.79 69.21 0.01 5,309 10
5 0.21 3.16 79.54 -0.01 5,444 7
Post-cutoff Period, 10/2023 to 12/2024
1 53.36 97.83 29.71 0.01 238 479
2 27.24 46.06 39.47 0.03 302 426
3 12.63 39.18 38.79 0.04 290 440
4 -6.04 23.52 37.14 -0.03 322 413
5 0.23 27.32 46.36 0.03 390 348
Panel B: With Context
Quintile MPE (%) MAPE (%) Directional Accuracy (%) Confidence Calibration Num Obs Refusals
Pre-cutoff, 01/1990 to 09/2023
1 1.35 10.80 74.46 0.01 18187 565
2 1.47 10.14 74.35 -0.01 18029 627
3 0.65 8.47 74.01 -0.03 17839 707
4 1.15 7.77 74.45 -0.01 17767 734
5 0.72 5.73 75.58 -0.06 17565 827
Recent Period, 01/2014 to 09/2023
1 1.86 10.69 73.35 -0.01 5376 135
2 0.97 8.55 73.62 -0.04 5311 163
3 0.45 6.68 72.88 -0.10 5192 231
4 0.47 5.48 74.52 -0.10 5083 236
5 0.35 3.29 80.56 -0.11 5217 234
Post-cutoff Period, 10/2023 to 12/2024
1 10.55 25.56 73.50 0.03 533 184
2 1.32 11.49 76.40 0.04 525 203
3 0.95 9.93 72.38 -0.11 456 274
4 0.90 8.74 72.18 -0.10 455 280
5 1.44 9.06 76.02 0.01 438 300
@@
The results show weaker recall compared to the Magnificent 7, with accuracy improving for larger stocks and recent periods. Without context in the pre-cutoff period, Mean Absolute Percent Errors are high: ranging from 11.29% for Quintile 5 containing the largest stocks to 24.70% for Quintile 1 containing the smallest stocks. Directional Accuracy is low, ranging from 42.66% to 61.11%. The high refusal rates suggest uncertainty for less prominent stocks. Over the recent 10-year period, errors decrease significantly ranging from 3.16% for Quintile 5 to 15.24% for Quintile 1. Directional Accuracy rises to 48.6%-79.5% and near-zero refusal rates, echoing the recency effect seen in macroeconomic levels. In the post-cutoff period, the error rates jump with Mean Absolute Percent Errors ranging from 23.52% to 97.83%. Directional Accuracy for all quintiles falls below 46.4%, with Quintile 1 having only 29.7% accuracy. With context, errors further improve for Directional Accuracy, with all quintiles reaching at least 74% in the pre-cutoff period. Larger stocks consistently show better recall, likely due to greater data prominence, aligning with the Magnificent 7’s stronger performance for high-profile securities.
@@
These findings complement our Magnificent 7 results, revealing that GPT-4o’s memorization weakens for less prominent stocks, particularly smaller ones, and is strongest for recent, larger-cap data. The high errors and refusals for small stocks contrast with the precision for META or GOOGL, suggesting memorization is skewed toward widely covered securities, similar to the robust recall of market indices and macroeconomic rates. The recency and context effect parallel improvements seen in prior tests, but the lower directional accuracy with context indicates limits in capturing trends for diverse portfolios. Again, this selective memorization underscores the risk of relying on LLMs for historical stock analysis, as apparent forecasting accuracy may stem from memorized prices of prominent stocks rather than broad predictive insight, reinforcing the need for post-cutoff evaluations.
Figure 7: Recall of exact numerical levels of closing prices by size quintiles.
This figure shows the LLM’s estimated closing prices for randomly selected stocks compared to the actual closing prices. Panels A, C, E, G, and I graph the actual values against the estimated values. Panels B, D, F, H, and J show the estimation error. Estimation error is calculated as (Estimated - Actual)/Actual and is shown in percentages (5 means 5%). The price plotted is the equal-weighted average price of the five quintile market cap groups, Q1 (smallest) through Q5 (largest). Each size quintile includes 50 randomly sampled stocks, resampled annually. The market cap breakpoints were calculated using NYSE stocks. The post-cutoff period (10/2023 onward) is shaded gray.
@@ 5.5 Can LLMs Follow “Fake” Knowledge Cutoff Prompts?
To assess whether GPT-4o can adhere to instructions not to use information beyond an artificially imposed cutoff date (i.e., fake cutoff), we test its performance in predicting U.S. real GDP growth rates over the period from March 1990 to June 2023.
@@
We design three prompting conditions using an artificial cutoff date of December 2010: (i) one where both system and user messages explicitly restrict knowledge to pre-2010 data, (ii) one where only the system prompt imposes this fake cutoff, and (iii) one where only the user prompt imposes this fake cutoff. The task requires the model to predict quarterly year-over-year GDP growth, with data split into pre-fake-cutoff (1990–2010) and post-fake-cutoff (2011 onward) periods to evaluate compliance with the constraint. Performance metrics include Mean Percent Error, Mean Absolute Percent Error, Directional Accuracy (correctly identifying up/down changes), and refusal counts, and are reported in Table 6.
@@
When both system and user prompts enforce the pre-2010 cutoff, GPT-4o performs well preceding this artificial cutoff, with a Mean Absolute Percent Error of 0.27% and Directional Accuracy of 95.12%, but struggles post-fake-cutoff, showing a higher error (0.59%) and lower accuracy (69.23%), with 38 refusals out of 51 post-fake-cutoff observations. By Proposition 1, this behavior is observationally equivalent to multiple contradictory explanations: the constraint genuinely working (the model cannot access post-2010 information), behavioral suppression (the model retains memorized information but learned to hide it), or strategic compliance (the model role-plays what it believes a restricted model would say). We also find that the results are similar when only the user prompt specifies the 2010 cutoff. In contrast, when only the system prompt specifies the cutoff, pre-fake-cutoff accuracy improves slightly (0.11% error, 97.59% accuracy). Still, post-fake-cutoff performance remains implausibly strong (0.05% error, 98.00% accuracy) with only one refusal across 50 observations. This minimal drop in accuracy post-fake-cutoff indicates that without reinforced user instructions, the model likely accesses memorized data, undermining the cutoff constraint.  [^11] We also test alternative pre-2005 and pre-2015 fake knowledge cutoffs for GDP growth in Appendix Tables A2 and A3, and find similar patterns.
In an additional set of analyses reported in Table A4, we conduct an analogous experiment for S&P 500 index levels at the daily frequency. In the Rolling Fake Cutoff test, we modify the user prompt to include instructions “Do not use any knowledge after time $\{t-1\}$” where $\{t-1\}$ is the date before the current date we ask the model to recall. We implement this procedure in the pre-cutoff period and the post-cutoff period and find the results to match the original recall procedure that does not include the fake knowledge cutoff instructions since the accuracy in the pre-cutoff period is still very high compared to the post-cutoff period. For the original benchmark, the model recalls pre-cutoff index levels with small errors (Mean Average Percent Error of 0.61% and Directional Accuracy of 80.6%), but performs poorly after the true cutoff (Mean Average Percent Error of 16.9% and Directional Accuracy of 45.7%). Under the Rolling Fake Cutoff instructions, the pre-cutoff performance remains strong (Mean Average Percent Error of 0.81% and Directional Accuracy of 69.6%), while the post-cutoff performance is weak (Mean Average Percent Error of 17.2% and Directional Accuracy of 42.2%). We also conduct two subsample tests for the periods 1990 to 2008 and 2009 to 2023 with two individual cutoff dates, December 31, 1999 and December 31, 2015, respectively. The results before and after the cutoff dates should be similar in both periods since all dates fall in the model’s pre-cutoff period. We do find this to be the case. For the 1990 to 2008 subsample, the Mean Average Percent Error actually falls from 1.39% pre-fake-cutoff to 0.67% post-fake-cutoff and Directional Accuracy rises from 59.3% to 74.4%. Similarly, in the 2009 to 2023 subsample, we find that the Mean Average Percent Error slightly falls from 0.05% to 0.03% and Directional Accuracy slightly rises from 97.3% to 98.1%.
These results align well with our earlier findings on macroeconomic indicators and market indices, where high pre-cutoff accuracy reflects memorization. The strong post-fake-cutoff performance without user prompt reinforcement (96% accuracy) demonstrates that prompt-based constraints empirically fail to prevent access to memorized data. Moreover, even if fake cutoffs had successfully reduced accuracy, Proposition 1 establishes we could not verify whether the constraint genuinely worked: reduced accuracy could reflect either successful restriction to historical information or the model learning to suppress known answers. These observationally equivalent outcomes make the target estimand fundamentally non-identified. The high refusal rate with dual prompts indicates partial compliance but not complete isolation from memorized information, illustrating the practical difficulty of constraining model outputs even when the theoretical impossibility of verification remains.
Table 6: Fake Knowledge Cutoff
@@
This table reports GPT-4o’s performance on U.S. real GDP growth predictions evaluated on data from the Philadelphia Fed’s Real-Time Data Set. We evaluate model accuracy under different artificially-imposed knowledge cutoff constraints: one where both system and user prompts reinforce the fake knowledge cutoff (pre-2010 only), another where only the system prompt specifies the cutoff, and one where only the user prompt specifies the cutoff. The task involves predicting the quarterly year-over-year GDP growth rate, with test data split into pre-fake-cutoff (1990–2010) and post-fake-cutoff (2011 onward) periods to assess whether the model respects the stated cutoff. Metrics include Mean Error (ME), Mean Absolute Error (MAE), Threshold Accuracy (percentage of guesses correctly above a threshold of 2.5%), Directional Accuracy (percentage of correct up/down changes), Confidence Calibration (correlation between the LLM’s confidence level and the MAPE), total observations, and refusal counts. The results indicate that explicitly instructing the model not to use post-2010 data yields higher refusal rates and weaker post-fake-cutoff performance, consistent with adherence to the knowledge constraint.
@@
ME (%) MAE (%) Threshold Accuracy (%) Directional Accuracy (%) Confidence Calibration Start Date End Date Num Obs Refusals
@@
Post fake-cutoff 0.58 0.59 69.23 75.00 0.51 03/01/2011 09/01/2023 13 38
@@
Post fake-cutoff -0.01 0.05 98.00 100.00 -0.24 03/01/2011 09/01/2023 50 1
@@
Post fake-cutoff 0.60 0.61 66.67 70.59 -0.15 03/01/2011 09/01/2023 18 32
@@ 5.6 Testing Masking Effectiveness
To evaluate whether masking can prevent GPT-4o from accessing memorized information, we examine its ability to identify the firm, year, and quarter of anonymized earnings conference call transcripts from January 2006 to September 2023, all within the pre-cutoff period before the model’s training knowledge cutoff of October 2023. Transcripts were anonymized using the entity neutering approach of Engelberg et al. (2025), removing identifying details like company names and dates. We focus on the Magnificent 7 stocks (AAPL, META, MSFT, GOOGL, NVDA, TSLA, AMZN) and portfolios grouped by market capitalization terciles (Large, Medium, Small), measuring performance through Mean Years Difference, Mean Absolute Years Difference, and accuracy in identifying the exact year, quarter and year, and firm. Results are reported in Table 7.
Table 7: Anonymized Conference Calls
This table reports GPT-4o’s performance in identifying the correct firm, year, and quarter of anonymized earnings conference call transcripts obtained from Capital IQ during the pre-cutoff period from January 2006 to September 2023. For each firm, we report the mean and mean absolute difference between the model’s recollection of the transcript years and the actual transcript years, as well as the percentage of calls for which the model correctly identified the exact year, the exact quarter and year, and the correct firm. Results are grouped by firm and by market capitalization terciles (Large, Mid, Small). Tercile breakpoints are formed using NYSE stocks.
@@
Pre-cutoff, 01/2006 to 09/2023
AAPL -0.03 0.05 95.95 93.24 100.00 74
META 0.22 0.27 73.47 2.04 100.00 49
MSFT -0.67 0.78 61.64 1.37 100.00 73
GOOGL -1.05 1.59 59.46 6.76 91.89 74
NVDA -1.34 1.66 50.00 1.61 93.55 62
TSLA -3.27 3.45 39.29 3.57 85.71 56
AMZN -1.87 2.11 30.67 4.00 85.33 75
@@
Pre-cutoff, 01/2006 to 09/2023
Large -1.54 1.92 44.69 8.09 65.77 593
Medium -1.87 2.37 39.93 7.34 40.78 586
Small -1.52 2.26 33.26 5.79 30.90 466
@@
GPT-4o demonstrates a remarkable ability to deanonymize transcripts, particularly for prominent firms. For AAPL, the model achieves 100% firm accuracy, 95.95% year accuracy, and 93.24% quarter-and-year accuracy, with a Mean Absolute Years Difference of just 0.05 years. META and MSFT also show perfect firm identification (100%), though year accuracy drops to 73.47% and 61.64%, respectively, and quarter-and-year accuracy is low (2.04% and 1.37%). Performance is slightly weaker for GOOGL (91.89% firm accuracy), NVDA (93.55%), TSLA (85.71%), and AMZN (85.33%), with year accuracy ranging from 30.67% to 59.46% and Mean Absolute Years Differences rising to 1.59–3.45 years. Across market cap terciles, firm accuracy declines from 65.77% for Large to 30.90% for Small stocks, with year accuracy dropping from 44.69% for Large to 33.26% for Small stocks, indicating stronger memorization for larger, more prominent firms, consistent with our portfolio stock findings.